06. High-altitude mountain climbers do not eat snow, but always melt To see why, calculate the energy absorbed from your body if you: t tirst with a stove (a) eat 1.0 kg of snow at 1.5 C which your body warms to body ten (b) melt 1.0 kg of snow -15° C using a stove and drink the resulting t 2°C, which your body has to warm to 37°C. 1,0 kg of water at An iron boiler of mass 180 kg contains 730 kg of water at 18°C. A heater sunl rate of 58000 kJ/h How long does it take for the water supplies energy at the (a) to reach the boiling point, and (b) to all have changed to steam? (II) The heat capacity, C, of an object is defined as the amount of heat needed to raise i temperature by 1 C° ts Thus, to raise the temperature by AT requires heat O given by Q CAT (a) Write the heat capacity C in terms of the specific heat, c, of the material. (b) What is the heat capacity of 1.0 kg of water? (c) Of 45 kg of water? (a) Estimate the total power radiated into space by the Sun, assuming it to be a perfect emer T 5500 K at The Suns radius is 7.0x 108 m (b) From this, determine the power per unit area arriving at the Earth, 1.5 x 10l m away
Answer #1

(6) (I) (a) Heat absorbed from body if you eat 1 kg of snow at -15°C which your body warms to 37°C

= M * Ci * (0 – (-15)) + Mi * L + M * Cw * (37 – 0 )

(M = Mass of water; Ci = specific heat of ice; L= latent heat of fusion; Cw = sp. Heat of water)

= 1*2108*15 + 1*333000 + 1*4187*37

= 519539 J or 519.539 KJ (Answer)

(b) Heat absorbed from body if you eat 1 kg of snow at -15°C (melted by stove up to 2° C) which your body warms to 37°C

= M * Cw * (37 – 2 )

= 1* 4187 * 35

= 146545 J or 146.545 KJ (Answer)

(7) (a) Heat capacity C in terms of specific heat c:

C = m * c where m = mass of substance   (Answer)

(b) Heat capacity of 1 kg of water = m * c = 1 kg * 4187 J/kg°C = 4187 J/°C (Answer)

(c) Heat capacity of 1 kg of water = m * c = 45 kg * 4187 J/kg°C = 188415 J/°C (Answer)

Thanks!!!